OpenStudy (anonymous):

Find the slope of the line with these given points (9, 3) and (3, -6).

OpenStudy (anonymous):

9/6 = 3/2

OpenStudy (anonymous):

slope of a line is y2-y1 / x2-x1 so, -6-3 / 3-9 = -9/-6 = 3/2

OpenStudy (anonymous):

Change in y over change in x... that means 3--6/9-3 9/6 3/2